I take 1g of zinc metal and after dissolving, diluting it to 500 ml and adjust the PH of 50 ml of it to make titration with an unknown concentration of EDTA.

I found 1g of zinc took 25 ml of EDTA that means 0.004 g of zinc equivalent to 1ml of unknown concentration of EDTA.

My question is how can I calculate the concentration of EDTA from the above information?



1
Expert's answer
2021-02-15T03:21:51-0500

1mL = 0.001L.

n=m/Mr = 0.004/65 = 6,15384615e-5 moles of Zn.

And Zn and EDTA equal.

Cm=n/V= 6,15384615e-5/0.001 = 0,0615384615 M.